# Ferromagnetism Mechanism in a Geometrically Frustrated Triangular Lattice

### v1

@@ Abstract
We study the emergent itinerant ferromagnetism and propose its underlying mechanism in the geometrically frustrated triangular lattice. Based on large-scale density matrix renormalization group simulations and unrestricted Hartree-Fock mean-field analysis, we identify itinerant ferromagnetic phases in the intermediate-$U$ Hubbard model with finite doping and reveal the kinetic mechanisms assisted by geometric frustration. Notably, we find that the doublon-singlon exchange process among other microscopic charge hoppings solely drives the fully polarized ferromagnetism for geometrically frustrated triangular lattice. Additionally, we establish the whole magnetic phase diagram and illustrate itinerant ferromagnetism within a finite range of electron doping for finite on-site Coulomb repulsion. The comparison of local spin correlations with recent cold-atom experiments is also discussed. Our work enhances the understanding of ferromagnetism mechanisms at intermediate coupling strength and finite doping concentrations.
Introduction.—It has been a long-standing issue in physics to understand the ferromagnetism mechanism in strongly correlated systems. As the fundamental theoretical model for diverse physics of these systems, the Hubbard model Tasaki (1998a, b); Vollhardt et al. (1999); Arovas et al. (2022); Qin et al. (2022) is initially formulated to understand the origin of ferromagnetism Gutzwiller (1963); Hubbard and Flowers (1963). The Hubbard Hamiltonian only includes the charge kinetic-energy term with the nearest-neighbor (NN) hopping amplitude $t$ and on-site Coulomb repulsion $U$,
@@
Given the lack of explicit or effective magnetic interactions favoring ferromagnetic order, ferromagnetism can only arise from the delicate interplay between kinetic energy and on-site repulsion. To date, rigorous theoretical results on the ferromagnetism mechanism in the Hubbard model have only been achieved in specific limits Tasaki (1998a, b); Vollhardt et al. (1999). The Stoner criterion Stoner (1938) states that ferromagnetism occurs when the product of the density of states at the Fermi energy $D_{\mathrm{F}}$ and on-site Coulomb repulsion $U$ satisfies the condition $UD_{\mathrm{F}}>1$. However, this criterion is derived from the Hartree-Fock approximation and is known to often overpredict ferromagnetic tendencies Vollhardt et al. (1996, 1999). Two other rigorous outcomes are Nagaoka ferromagnetism Thouless (1965); Nagaoka (1966), arising from kinetic energy minimization of a single doped hole at $U=\infty$, and flat-band ferromagnetism Mielke (1991a); Tasaki (1992) with dispersionless lowest energy bands. Over the past several decades, it has attracted substantial efforts to explore ferromagnetism beyond those well-known limits Riera and Young (1989); Basile and Elser (1990); Emery et al. (1990); Shastry et al. (1990); Mielke (1991b, 1992); Hanisch et al. (1995); Wurth et al. (1996); Brunner and Muramatsu (1998); Ulmke (1998); Becca and Sorella (2001); Merino et al. (2006); Chang et al. (2010, 2011); Carleo et al. (2011); Liu et al. (2012); Schlömer et al. (2023); Davydova et al. (2023); Samajdar and Bhatt (2023); He et al. (2023); Morera and Demler (2024), particularly for the square lattices Riera and Young (1989); Basile and Elser (1990); Emery et al. (1990); Shastry et al. (1990); Wurth et al. (1996); Brunner and Muramatsu (1998); Carleo et al. (2011); Liu et al. (2012). However, revealing the ferromagnetism mechanism in the ground state of geometrically frustrated triangular lattices at intermediate $U$ and finite doping remains challenging Merino et al. (2006); Samajdar and Bhatt (2023); He et al. (2023); Morera and Demler (2024).
More recently, the realization of cold-atom quantum simulators of the triangular-lattice Hubbard model in optical lattices Yang et al. (2021); Xu et al. (2023); Lebrat et al. (2024); Prichard et al. (2024) has significantly advanced this issue. With unprecedented control over doping and coupling strength Chin et al. (2010); Lewenstein et al. (2012), the ferromagnetic correlations are directly observed Xu et al. (2023); Lebrat et al. (2024); Prichard et al. (2024) at intermediate on-site Coulomb repulsion and finite particle doping. In a different context, the transition metal dichalcogenides (TMD) moiré materials provide another avenue for simulating the triangular-lattice Hubbard model, wherein ferromagnetism has also been observed Tang et al. (2020); Ciorciaro et al. (2023); Anderson et al. (2023); Seifert and Balents (2024). These experimental advancements necessitate theoretical study of the observed ferromagnetism and its microscopic mechanisms in the geometrically frustrated triangular lattice. Unlike the intensively studied hole-doped case, experimentally observed ferromagnetism occurs at the electron-doped side, the physics of which could be quite different due to the particle-hole asymmetry Zhu and Chen (2023); Chen et al. (2022); Xu et al. (2023); Prichard et al. (2024). In the limit of nearly half filling $n$$\to$1111 ($n$ denotes electron density per site) and full filling $n$$\to$2222, the Nagaoka-type Hanisch et al. (1995); Chen et al. (2022); Morera et al. (2023); Davydova et al. (2023); Schlömer et al. (2023); Lee et al. (2023); Morera and Demler (2024) and Müller-Hartmann-type ferromagnetism Li et al. (2023) have been proposed, respectively. However, the ferromagnetism mechanism for intermediate density around $n=3/2$ is still an open issue. At $n=3/2$, the Fermi surface nesting and van Hove singularities exist at $U=0$, weak interactions would trigger instabilities towards magnetic orders with nesting wave vectors Martin and Batista (2008); Akagi and Motome (2010). However, when the weak-coupling theory fails at the strong-coupling regime Pasrija and Kumar (2016), it is still a much-needed endeavor to identify the magnetism and its mechanism.
Motivated by the above, we explore the ferromagnetism and its underlying mechanism in the triangular-lattice Hubbard model from intermediate to infinite interactions near electron doping $\delta=1/2$ (i.e., $n=3/2$, where the doping concentration is defined by $\delta\equiv n-1$). Using large-scale density matrix renormalization group (DMRG) and unrestricted Hartree-Fock (UHF) methods, we first present the emergence of ferromagnetic ground state with metallic charge behavior at $n=3/2$ as the on-site Coulomb repulsion $U$ increases from intermediate to strong and ultimately infinite values. By examining distinct microscopic hopping processes inherent in the Hubbard model at intermediate $U$, we reveal that the doublon-singlon exchange, assisted by the geometric frustration inherited from the lattice, solely contributes to the fully polarized ferromagnetism. We further confirm our conclusions by selectively deactivating each hopping process and verifying their individual impacts on itinerant ferromagnetism. Additionally, we demonstrate that itinerant ferromagnetism around $\delta=1/2$ persists within a specific range of electron doping and find the local spin correlations at zero temperature resemble those observed in a recent cold-atom experiment Xu et al. (2023).
Fig. 1: Evolution of the magnetic and electronic structure across interaction $U/t$ at electron doping $\delta=1/2$. (a) Ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})$ as a function of $U/t$. (b) Polarization $P$ from unrestricted Hartree-Fock method with periodic boundary conditions. (c) Line-cut plot of the static spin structure factor $S(\mathbf{q})$ along momentum path through $\mathbf{\Gamma}$ point (white dashed line in inset). Labels are for different interactions $U/t$. (d) Single-particle propagator $|C(r)|$ for different $U/t$, compared to $\sim r^{-2}$ and $\sim r^{-1}$. (e) Line-cut plot of electron momentum distribution $n(\mathbf{k})$ along momentum path through the momentum $\mathbf{\Gamma}$ (blue dashed line in inset). Insets: contour plots of $S(\mathbf{q})$ in (c) and $n(\mathbf{k})$ in (e) with $U=24$, $N=24\times 6$. White hexagon: the first Brillouin zone.
Fig. 2: Microscopic hopping processes in the Hubbard model and their kinetic energy contributions for intermediate $U/t$. (a)-(c) Schematic illustration of nearest-neighbor hopping processes between (a) a doubly occupied site and an empty site, or two singly occupied sites; (b) a singly occupied site and an empty site; (c) a doubly occupied site and a singly occupied site. These processes are described in (3). (d) Schematic illustration of ring-exchange coupling. (e),(f) The kinetic energies (in blue) of different microscopic hopping processes as functions of $U/t$ for $L_{y}=4$ in (e) and $L_{y}=6$ in (f). Ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})$ (in orange) is plotted for comparison.
Itinerant Ferromagnetism at Electron Doping $\delta=1/2$.— We first identify the nature of the ground state in the triangular-lattice Hubbard model at electron doping $\delta=1/2$ using DMRG and UHF (details see supplementary SM ).
To examine the magnetic properties of the ground state, we use DMRG to compute the static spin structure factor $S(\mathbf{q})$, defined by $S(\mathbf{q})\equiv\sum_{\mathbf{i},\mathbf{j}}\langle{\mathbf{S}_{\mathbf{i}}\cdot\mathbf{S}_{\mathbf{j}}}\rangle e^{i\mathbf{q}\cdot(\mathbf{i}-\mathbf{j})}/N$. Magnetic order is signaled by peaks at certain wave vectors $\mathbf{q}=\mathbf{q}_{0}$. The DMRG simulations are performed on the cylinder with size $N=L_{x}\times L_{y}$, where $L_{x}$ ($L_{y}$) represents the length of the cylinder (circumference). We mainly study $L_{y}=4$, $6$ and push bond dimension $D$ up to $D\approx 36,000$. Notably, large bond dimensions are necessary for distinguishing ferromagnetic and incommensurate spin-density-wave (SDW) phases (see supplementary SM ). We first focus on the evolution of the momenta $\mathbf{q}_{0}$ as a function of interaction strength $U/t$. Our DMRG results reveal that, with increasing $U/t$ , $\mathbf{q}_{0}$ progressively shifts from $\mathbf{M}$ to $\mathbf{\Gamma}$ points SM and finally fixed at the $\mathbf{\Gamma}$ beyond a critical strength $U_{c}/t$. We therefore keep track of the squared order parameter $m_{s}^{2}(\mathbf{\Gamma})\equiv S(\mathbf{\Gamma})/N$ as a function of $U/t$. As shown in Figs. 1(a), (c), when further increasing $U/t$, $m_{s}^{2}(\mathbf{\Gamma})$ gradually increases until plateauing, signaling the establishment of a stabilized ferromagnetic phase at larger $U/t$ and the evolution from partially to fully polarized phases. The saturated value of $m_{s}^{2}(\mathbf{\Gamma})$ matches the theoretically expected $\delta_{\mathbf{q},\mathbf{\Gamma}}/16$ at $n=3/2$.
To confirm the robustness of ferromagnetism in the thermodynamic limit, we further employ the unrestricted Hartree-Fock (UHF) mean-field method SM . In the UHF method, under an uncorrelated state ansatz, both the on-site densities $\langle n_{\mathbf{i},\sigma}\rangle$ and the spin-flips $\langle S_{\mathbf{i}}^{-}\rangle$ and $\langle S_{\mathbf{i}}^{+}\rangle$ are involved to factorize the Hubbard interaction
$n_{\mathbf{i},\uparrow}n_{\mathbf{i},\downarrow}\rightarrow$ $\langle n_{\mathbf{i},\uparrow}\rangle n_{\mathbf{i},\downarrow}+\langle n_{\mathbf{i},\downarrow}\rangle n_{\mathbf{i},\uparrow}-\langle n_{\mathbf{i},\uparrow}\rangle\langle n_{\mathbf{i},\downarrow}\rangle$
$-\langle S_{\mathbf{i}}^{+}\rangle S_{\mathbf{i}}^{-}-\langle S_{\mathbf{i}}^{-}\rangle S_{\mathbf{i}}^{+}+\langle S_{\mathbf{i}}^{-}\rangle\langle S_{\mathbf{i}}^{+}\rangle.$ (2)
Here $S_{\mathbf{i}}^{+}=c_{\mathbf{i},\uparrow}^{\dagger}c_{\mathbf{i}\downarrow}$ and $S_{\mathbf{i}}^{-}=c_{\mathbf{i},\downarrow}^{\dagger}c_{\mathbf{i},\uparrow}$ are spin ladder operators. We have a total of $4N$ variational parameters, obtained through iterative diagonalization of the Hartree-Fock Hamiltonian. To ensure stable convergence, we utilize direct inversion in the iterative subspace method Pulay (1980), and the convergence criterion for the energy is set to $10^{-8}t$. As shown in Fig. 1(b), the polarization $P\equiv 2Nm_{s}(\mathbf{\Gamma})$ increases gradually as $U/t$ becomes larger, and finally a full polarization state occurs at $U\simeq 4.5t$ independently of system size. This aligns with Fig. 1(a) and is also consistent with the recent Hartree-Fock analysis He et al. (2023), indicating robust fully-polarized ferromagnetism in the strong coupling regime.
@@
Having established the ferromagnetism in the spin channel, we further examine the electronic properties in the charge channel. We compute the single-particle propagator $C(r)\equiv\sum_{\sigma}\langle c_{\mathbf{i}_{0},\sigma}^{\dagger}c_{{\mathbf{i}_{0}+r\bm{e}_{x}},\sigma}\rangle$, and the momentum distribution of the electrons $n(\mathbf{k})\equiv\sum_{\mathbf{i},\mathbf{j},\sigma}\langle c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}\rangle e^{i\mathbf{k}\cdot(\mathbf{i}-\mathbf{j})}/N$. Figure 1(d) illustrates the power-law decay of $|C(r)|$ with an exponent approximately equal to 1, identifying the gapless nature of the electrons in the ferromagnetic phase. In the inset of Fig. 1(e), the contour plot of $n(\mathbf{k})$ exhibits an abrupt change from 2 to 1 at certain momenta, as more clearly illustrated by the line-cut plot of $n(\mathbf{k})$ through the momentum $\mathbf{\Gamma}$ in the main panel of Fig. 1(e). This abrupt change indicates the presence of a well-defined Fermi surface, with the position of this sudden change characterizing the Fermi momenta. These observations are robust for the strength of intermediate-to-infinite interactions $U/t$ and different system sizes $N$, suggesting that the electrons in the ferromagnetic phase for both intermediate $U/t$ and infinite $U/t$ limit are itinerant.

### v2

# Geometric Frustration Assisted Kinetic Ferromagnetism in Doped Mott Insulators

@@ Abstract
Understanding ferromagnetism mechanism in doped Mott insulators on frustrated lattices remains challenging at intermediate coupling and finite doping. Here, we study the itinerant ferromagnetism and propose its mechanism in doped Mott insulators on a geometrically frustrated triangular lattice. Using large-scale density matrix renormalization group (DMRG) and unrestricted Hartree-Fock mean-field methods, we reveal that itinerant ferromagnetism appears at intermediate coupling ($10\lesssim U\ll\infty$) near 50% electron doping in the triangular-lattice Hubbard model. By analyzing all microscopic hopping processes, we find that doublon-singlon exchange alone drives the fully polarized ferromagnetism and uncovers the particle-hole asymmetry. We also establish the magnetic phase diagram and compare local spin correlations with recent experiments. Random phase approximation and DMRG calculations consistently confirm that the ferromagnetism persists when $SU(2)$ symmetry is explicitly broken by magnetic anisotropy. These results clarify a microscopic route to itinerant ferromagnetism at intermediate coupling and finite doping in doped Mott insulators.
Introduction.
It has been a long-standing issue in physics to understand the ferromagnetism mechanism in strongly correlated systems. As the fundamental theoretical model for diverse physics of these systems, the Hubbard model is initially formulated to understand the origin of ferromagnetism Gutzwiller1963 ; Hubbard1963 . Given the lack of explicit or effective magnetic interactions favoring ferromagnetic order, ferromagnetism can only arise from the delicate interplay between kinetic energy and on-site repulsion $U$. To date, rigorous theoretical results on the ferromagnetism mechanism in the Hubbard model have only been achieved in specific limits Tasaki1998 ; Tasaki1998Review ; Vollhardt1999 . The Stoner criterion Stoner1938 states that ferromagnetism occurs when the product of the density of states at the Fermi energy $D_{\mathrm{F}}$ and on-site Coulomb repulsion $U$ satisfies the condition $UD_{\mathrm{F}}>1$. However, this criterion is derived from the Hartree-Fock approximation and is known to often overpredict ferromagnetic tendencies Vollhardt1996 ; Vollhardt1999 . Two rigorous outcomes are Nagaoka ferromagnetism Thouless1965 ; Nagaoka1966 , arising from kinetic energy minimization of a single doped charge at $U=\infty$, and flat-band ferromagnetism Mielke1991b ; Tasaki1992 with dispersionless lowest energy bands. Over the past several decades, it has attracted substantial efforts to explore ferromagnetism beyond those well-known limits Riera1989 ; Basile1990 ; Emery1990 ; Shastry1990 ; Mielke1991 ; Mielke1992 ; Hanisch1995 ; Wurth1996 ; Brunner1998 ; Ulmke1998 ; Becca2001 ; Merino2006 ; ChiaChenChang2010 ; SoonYongChang2011 ; Carleo2011 ; LiLiu2012 ; GangLi2014 ; Schlomer2023 ; Davydova2023 ; Samajdar2023 ; YuchiHe2023 ; Morera2024 ; Lorenzo2024 , particularly for the square lattices Riera1989 ; Basile1990 ; Emery1990 ; Shastry1990 ; Wurth1996 ; Brunner1998 ; Carleo2011 ; LiLiu2012 . However, understanding the microscopic mechanism for ferromagnetism in geometrically frustrated triangular lattices at both intermediate Coulomb repulsion ($10\lesssim U\ll\infty$) and finite doping $\delta$ remains an open issue with fundamental challenge Merino2006 ; Samajdar2023 ; YuchiHe2023 ; Morera2024 ; Samajdar2024PRB .
More recently, the transition metal dichalcogenides (TMD) moiré materials YanhaoTang2020 ; Ciorciaro2023 ; Anderson2023 ; Seifert2024 and the quantum simulators in optical lattices YangJin2021 ; MuqingXu2023 ; Lebrat2024 ; Prichard2023 ; ChinCheng2010 ; Lewenstein2012 have significantly advanced this issue. The essential physics of these experimental platforms could be captured by the triangular-lattice Hubbard model. With unprecedented control over doping and coupling strength, ferromagnetism has been observed at an intermediate coupling and large electron doping. These experimental advancements also necessitate theoretical study on the mechanisms and stability of the observed ferromagnetism in triangular lattices.
Unlike the intensively studied hole-doped case, experimentally observed ferromagnetism occurs at the electron-doped side, the physics of which could be quite different due to the particle-hole asymmetry Zhu2023SC ; ShuaiChen2022SCBA ; MuqingXu2023 ; Prichard2023 ; GuanhuaHuang2023 . In the limit of nearly half filling $n\to$11 ($n$ denotes electron density per site) and full filling $n\to$22, the Nagaoka-type Hanisch1995 ; ShuaiChen2022SCBA ; Morera2023 ; Davydova2023 ; Schlomer2023 ; KyungminLee2023 ; Morera2024 and Müller-Hartmann-type ferromagnetism ChengshuLi2023 have been proposed, respectively. However, the ferromagnetism mechanism for intermediate density around $n=3/2$, which is significantly away from half-filling and full-filling, remains an open issue. At $n=3/2$, weak interactions would trigger instabilities towards magnetic orders with nesting wave vectors Martin2008 ; Akagi2010 . However, when the weak-coupling theory fails at the strong-coupling regime Pasrija2016 , it is still a much-needed endeavor to identify the magnetism and its mechanism, particularly for the intermediate coupling regions.
Fig. 1: $|$ The static spin structure factor $\boldsymbol{S(\mathbf{q})}$ as a function of coupling strength $\boldsymbol{U/t}$ at electron doping $\boldsymbol{\delta=1/2}$. (a)-(b) The contour plot of $S(\mathbf{q})$ for cylinders at $U/t=10$ (a), and $U/t=24$ (b). The black dots represent the accessible momenta in the Brillouin zone (denoted as a white hexagon), and interpolation has been applied in the contour plot. We remark that the cylindrical geometry adopted in the density matrix renormalization group (DMRG) simulation breaks the lattice’s rotational symmetry. (c) Schematic of the evolution of the $\mathbf{q}_{0}$ with increasing $U/t$, indicated by the color gradient from light to dark along the black dashed arrows. Here, $\mathbf{q}_{0}$ denotes the momenta where $S(\mathbf{q})$ reaches its maximum, within the Brillouin zone represented by a hexagon. Rotational symmetry of triangular lattices is assumed in this schematic for clarity. (d) Ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})$ as a function of $U/t$. For larger $U/t$, the peak of $S(\mathbf{q})$ stabilizes at momentum $\mathbf{\Gamma}$ with a size-independent saturation value, indicating the fully polarized ferromagnetism. (e) Line-cut plot of $S(\mathbf{q})$ along the momentum path passing through the $\mathbf{\Gamma}$ point in (c), depicted as a white dashed line in (b). (f) Squared order parameter $m_{s}^{2}(\mathbf{\Gamma})$ from unrestricted Hartree-Fock calculations with periodic boundary conditions.
Fig. 2: $|$ The single-particle propagator $\boldsymbol{C(r)}$ at separation $\boldsymbol{r}$ and electron momentum distribution $\boldsymbol{n(\mathbf{k})}$ at electron doping $\boldsymbol{\delta=1/2}$. (a) Single-particle propagator $|C(r)|$ for different $U/t$, where $U$ is the on-site Hubbard interaction, and $t$ is the nearest-neighbor hopping amplitude. For comparison, the dashed and solid gray lines correspond to $|C(r)|\sim r^{-2}$ and $|C(r)|\sim r^{-1}$, respectively. (b) Line-cut plot of electron momentum distribution $n(\mathbf{k})$ along a specific momentum path that traverses the momentum $\mathbf{\Gamma}$, depicted as a blue dashed line in its inset. The inset shows the contour plot of $n(\mathbf{k})$, the white hexagon denotes the first Brillouin zone, and the colour bar encodes $n(\mathbf{k})$.
Motivated by the above, we explore the ferromagnetism and its underlying mechanism in the triangular-lattice Hubbard model from intermediate to infinite interactions near electron doping $\delta=1/2$ (i.e., $n=3/2$, where the doping concentration is defined by $\delta\equiv n-1$). Using large-scale density matrix renormalization group (DMRG) and unrestricted Hartree-Fock (UHF) methods, we first present the emergence of ferromagnetic ground state with metallic charge behavior at $n=3/2$ as the on-site Coulomb repulsion $U$ increases from intermediate to strong and ultimately infinite values. Among distinct microscopic hopping processes inherent in a decomposition of the intermediate-$U$ Hubbard model, we reveal that only the doublon-singlon exchange—assisted by the geometric frustration of the lattice—contributes to the fully polarized ferromagnetism. We further offer stronger evidence for this conclusion by selectively deactivating each hopping process and verifying their individual impacts on itinerant ferromagnetism. Additionally, we demonstrate that itinerant ferromagnetism around $\delta=1/2$ persists within a specific range of electron doping and find the local spin correlations at zero temperature resemble those observed in a recent experiment MuqingXu2023 . Finally, we confirm that ferromagnetism is robust against the explicit breaking of $SU(2)$ symmetry in the Hubbard model through the introduction of anisotropy.
RESULTS
Itinerant Ferromagnetism at Intermediate $U$ and Large Electron Doping. In this work, we study the ground-state properties of the triangular-lattice Hubbard model at intermediate coupling strength. The many-body Hamiltonian reads
@@ RESULTS
Here, $t>0$ and $U>0$ are the nearest-neighbor (NN) hopping amplitude and the on-site Coulomb repulsion energies. The operators $c_{\mathbf{i},\sigma}^{(\dagger)}$ ($c_{\mathbf{i},\sigma}$) creates (annihilates) an electron on site $\mathbf{i}$ with spin $\sigma=\uparrow,\downarrow$, and $n_{\mathbf{i},\sigma}\equiv c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{i},\sigma}$ is the number operator. The summation of bonds $\langle\cdots\rangle$ runs over all NN bonds in the triangular lattice. The doping concentration is represented by $\delta=n-1$, with $n$ indicating the electron density per site. Here, $\delta=0$ corresponds to the half-filling, and $\delta>0$ ($\delta<0$) is the electron-doped (hole-doped) case. Here, we fix $t=1$.
We first identify the itinerant ferromagnetic phase at electron doping $\delta=1/2$ and intermediate coupling strength in the triangular-lattice Hubbard model using DMRG and UHF methods. Unless stated otherwise, the figures in the following show DMRG data obtained on $L_{x}\times L_{y}$ cylindrical lattices with primitive vectors $\boldsymbol{e}_{x}=(1,0)$ and $\boldsymbol{e}_{y}=(1/2,\sqrt{3}/2)$, open boundary conditions along $L_{x}$ (length), and periodic along $L_{y}$ (circumference). UHF is applied with periodic boundary conditions in both directions. We remark that our largest DMRG simulations are performed on systems with width up to ${L_{y}}=6$ and ${L_{x}}\leq 24$. These system sizes (i) significantly suppress finite-size effects and (ii) are crucial for clearly distinguishing ferromagnetism from the incommensurate spin density wave (iSDW) correlations.
To identify the nature of the ground state in the spin and charge channels at electron doping $\delta=1/2$, we examine the evolution of the magnetic and electronic structure across a range of interaction strengths, from finite to infinite $U/t$, as depicted in Fig. 1 and Fig. 2. The magnetic order is signaled by peaks at certain wave vectors $\mathbf{q}=\mathbf{q}_{0}$ in the static spin structure factor $S(\mathbf{q})$, defined by $S(\mathbf{q})\equiv\sum_{\mathbf{i},\mathbf{j}}\langle{\mathbf{S}_{\mathbf{i}}\cdot\mathbf{S}_{\mathbf{j}}}\rangle e^{i\mathbf{q}\cdot(\mathbf{i}-\mathbf{j})}/N$.
We first focus on the evolution of the momenta $\mathbf{q}_{0}$ as a function of interaction strength $U/t$. In the noninteracting limit ($U/t=0$), electron doping $\delta=1/2$ corresponds to the Van-Hove filling at which a perfect nesting of the Fermi surface exists. At weak coupling side $U/t\to 0$, it is expected that nesting leads to peaks of $S(\mathbf{q})$ locating at the $\mathbf{M}$ points in the Brillouin zone [as labeled in Fig. 1c]. Previous weak-coupling theory Martin2008 proposes that the candidate ground state exhibits noncoplanar chiral magnetic order. When the weak coupling assumption is no longer applicable at intermediate to strong interaction strengths $U/t$, using DMRG we numerically observe that other magnetic ordering vectors distinct from the nesting vectors emerge, as shown in Figs. 1a. Interestingly, as $U/t$ increases, the locations of these peaks in $S(\mathbf{q})$ shift towards the $\mathbf{\Gamma}$ point [as labeled in Fig. 1c], which is also consistent with the recent Hartree-Fock analysis YuchiHe2023 . Upon surpassing a critical strength $U_{c}/t$ toward the limit $U/t=\infty$, the peak of $S(\mathbf{q})$ becomes fixed at $\mathbf{\Gamma}$ point, as illustrated in Fig. 1b, d, indicating the establishment of a stabilized ferromagnetic phase in the strong coupling regime. In Fig. 1c, we schematically plot such evolution of the peaks in $S(\mathbf{q})$ with increasing $U/t$, as visualized by the dashed arrow from momenta $\mathbf{M}$ to $\mathbf{\Gamma}$, and the additional supporting numerical data are provided in the Supplementary Note 1. Notably, after the peak of $S(\mathbf{q})$ shifts to the $\mathbf{\Gamma}$ point, its height gets larger in a finite range of increasing $U/t$ [see Fig. 1e], until reaching its maximum value. This signals the evolution from a partially-polarized to a fully-polarized phase. We therefore keep track of the squared order parameter $m_{s}^{2}(\mathbf{\Gamma})\equiv S(\mathbf{\Gamma})/N$ as a function of $U/t$. As illustrated by Fig. 1d, when increasing $U/t$, $m_{s}^{2}(\mathbf{\Gamma})$ gradually increases until plateauing, signaling the establishment of a stabilized ferromagnetic phase at intermediate $U/t$. The saturated value matches the ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})\sim\delta_{\mathbf{q},\mathbf{\Gamma}}/16$ for a fully polarized state at electron doping $\delta=1/2$.
@@
Having established the ferromagnetism in the spin channel, we further examine the electronic properties in the charge channel. We compute the single-particle propagator $C(r)\equiv\sum_{\sigma}\langle c_{\mathbf{i}_{0},\sigma}^{\dagger}c_{{\mathbf{i}_{0}+r\boldsymbol{e}_{x}},\sigma}\rangle$, and the momentum distribution of the electrons $n(\mathbf{k})\equiv\sum_{\mathbf{i},\mathbf{j},\sigma}\langle c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}\rangle e^{i\mathbf{k}\cdot(\mathbf{i}-\mathbf{j})}/N$. Fig. 2a illustrates the power-law decay of $|C(r)|$ with an exponent approximately equal to 1, identifying the gapless nature of the electrons in the ferromagnetic phase. In the inset of Fig. 2b, the contour plot of $n(\mathbf{k})$ exhibits an abrupt change from 2 to 1 at certain momenta, as more clearly illustrated by the line-cut plot of $n(\mathbf{k})$ through the momentum $\mathbf{\Gamma}$ in the main panel of Fig. 2b. This abrupt change indicates the presence of a well-defined Fermi surface, with the position of this sudden change characterizing the Fermi momenta. These observations are robust for the strength of intermediate-to-infinite interactions $U/t$ and different system sizes $N$, suggesting that the electrons in the ferromagnetic phase for both intermediate $U/t$ and infinite $U/t$ limit are itinerant.
To confirm the robustness of ferromagnetism in the thermodynamic limit, we further employ the unrestricted Hartree-Fock (UHF) mean-field calculations. In the UHF method, under an uncorrelated state ansatz, both the on-site densities $\langle n_{\mathbf{i},\sigma}\rangle$ and the spin-flips $\langle S_{\mathbf{i}}^{-}\rangle$ and $\langle S_{\mathbf{i}}^{+}\rangle$ are involved to factorize the Hubbard interaction. Here $S_{\mathbf{i}}^{+}=c_{\mathbf{i},\uparrow}^{\dagger}c_{\mathbf{i}\downarrow}$ and $S_{\mathbf{i}}^{-}=c_{\mathbf{i},\downarrow}^{\dagger}c_{\mathbf{i},\uparrow}$ are spin ladder operators. Then the Hartree-Fock factorizes the Hamiltonian to
$H_{\mathrm{HF}}=$ $-\sum_{\langle\mathbf{ij}\rangle}(t+\mu\delta_{\mathbf{ij}})c_{\mathbf{i},\sigma}^{\dagger}c_{\mathbf{j},\sigma}$
$+\sum_{\mathbf{i}}U[\langle n_{\mathbf{i},\uparrow}\rangle n_{\mathbf{i},\downarrow}+\langle n_{\mathbf{i},\downarrow}\rangle n_{\mathbf{i},\uparrow}-\langle S_{\mathbf{i}}^{+}\rangle S_{\mathbf{i}}^{-}-\langle S_{\mathbf{i}}^{-}\rangle S_{\mathbf{i}}^{+}]$
$+\sum_{\mathbf{i}}U\left[-\langle n_{\mathbf{i},\uparrow}\rangle\langle n_{\mathbf{i},\downarrow}\rangle+\langle S_{\mathbf{i}}^{-}\rangle\langle S_{\mathbf{i}}^{+}\rangle\right],$ (2)
where the chemical potential $\mu$ controls the electron doping $\delta=1/2$. The mean-field parameters are expectation values under the ground state $|\Omega\rangle$ of the $H_{\mathrm{HF}}$,
$\langle n_{\mathbf{i},\sigma}\rangle\equiv\langle\Omega|n_{\mathbf{i},\sigma}|\Omega\rangle,\quad\langle S_{\mathbf{i}}^{\pm}\rangle\equiv\langle\Omega|S_{\mathbf{i}}^{\pm}|\Omega\rangle,$ (3)
and the on-site densities are subject to the particle number constraint $\sum_{\mathbf{i}}[\langle n_{\mathbf{i},\uparrow}\rangle+\langle n_{\mathbf{i},\downarrow}\rangle]=N(1+\delta)$ with $N$ being the number of sites. There are a total $4N$ variational parameters, which can be iteratively solved by the self-consistent equations
$\langle n_{\mathbf{i},\sigma}\rangle$ $=\sum_{n}f_{FD}(E_{n})\langle\psi_{n}|n_{\mathbf{i},\sigma}|\psi_{n}\rangle$ (4)
$\langle S_{\mathbf{i}}^{\pm}\rangle$ $=\sum_{n}f_{FD}(E_{n})\langle\psi_{n}|S_{\mathbf{i}}^{\pm}|\psi_{n}\rangle$ (5)
where $|\psi_{n}\rangle$ is an eigenstate of $H_{\mathrm{HF}}$ with $H_{\mathrm{HF}}|\psi_{n}\rangle=E_{n}|\psi_{n}\rangle$ and $f_{FD}(E_{n})$ is the Fermi-Dirac distribution. For a stable convergence, we utilize the direct inversion in the iterative subspace (DIIS) method Pulay1980 , and the convergence criterion for the energy is set to $10^{-8}t$.
As shown in Fig. 1f, the squared order parameter $m_{s}^{2}(\mathbf{\Gamma})$ increases gradually as $U/t$ becomes larger, and finally a full polarization state occurs at $U\simeq 4.5t$ independently of system size. This aligns with Fig. 1d and is also consistent with the recent Hartree-Fock analysis YuchiHe2023 , indicating robust fully-polarized ferromagnetism in the strong coupling regime.
Fig. 3: $|$ Microscopic hopping processes in the Hubbard model and their kinetic energy contributions for intermediate coupling $\boldsymbol{U/t}$. (a)-(c) Schematic illustration of nearest-neighbor hopping processes described in Eq. (6). (d) Schematic illustration of ring-exchange coupling. (e),(f) The kinetic energies (in blue) of different microscopic hopping processes as functions of $U/t$ for lattice width $L_{y}=4$ in (e) and $L_{y}=6$ in (f). Ferromagnetic squared order parameter $m_{s}^{2}(\mathbf{\Gamma})$ (in orange) is plotted for comparison.